# Multi-agent Embodied AI: Advances and Future Directions

### v1

@@ 2 Preliminaries
In this section, we present the core techniques that underpin Embodied AI, starting with a formal definition of Embodied AI itself. We also define the concept of MAS and explore a range of essential methods, including optimal control (OC), Reinforcement Learning (RL), MARL, Hierarchical Learning, IL, and Generative Models.
@@ 2.1 Embodied AI
Embodied AI ( Figure 2) refers to a class of intelligent agents equipped with physical bodies that enable them to perceive, act upon, and adapt to their environments through continuous interaction [109]. The conceptual roots of embodied AI trace back to Alan Turing’s early proposition in the 1950s, suggesting that genuine intelligence must arise from sensory and motor experiences rather than purely symbolic computation [179]. This notion was further formalized in the 1980s through embodied cognition theory, which posits that cognition is inherently shaped by the agent’s physical form and interaction with the world [19]. In contrast to traditional AI paradigms that rely on abstract reasoning or passive learning from static datasets, embodied AI emphasizes real-world interaction as the foundation for learning and decision-making.
At a system level, embodied AI architectures are typically composed of three tightly integrated components: perception, cognition, and action. Agents use physical sensors to acquire real-time, context-dependent information from their surroundings. This sensory data is processed by cognitive modules that support reasoning, interpretation, and planning. Resulting decisions are then translated into physical actions via actuators, which modify the environment and initiate new perceptual inputs. These processes form a continuous feedback loop known as the perception–cognition–action cycle [16], which enables embodied agents to adapt their behavior dynamically based on environmental feedback. Central to the embodied AI paradigm are three foundational attributes that govern how intelligence emerges and evolves within physical agents:
  • Embodiment Embodied AI is rooted in a physical agent equipped with the capabilities to perceive, move, and act in the real world. These agents come in diverse forms, including humanoid robots, quadrupeds, autonomous ground vehicles, and aerial drones. The physical body serves not only as the medium through which the agent interacts with its environment, but also as the structural foundation that constrains and enables its behavioral repertoire. The morphology, sensorimotor fidelity, and actuation of the body jointly define the range and granularity of the agent’s possible interactions, shaping the scope of its situated intelligence.
@@
Recent advancements in generative models, particularly LLMs [251], have further expanded the cognitive capabilities of embodied agents. By leveraging their powerful reasoning and generalization abilities, LLMs enable embodied systems to understand language instructions, ground semantic knowledge in physical experience, and perform zero-shot or few-shot adaptation. These developments have accelerated the deployment of embodied AI in real-world domains such as robotics, autonomous driving, intelligent manufacturing, and healthcare [115]. Importantly, embodied AI is not merely the integration of powerful AI models with robotic platforms; rather, it represents a co-evolutionary paradigm where intelligent algorithms (“brains”), physical structures (“bodies”), and dynamic environments evolve jointly to support adaptive, embodied intelligence.
@@ 2.2 Multi-Agent Systems
MAS are composed of multiple autonomous agents, each capable of perceiving its environment, making independent decisions, and executing actions accordingly [37]. In contrast to traditional centralized control paradigms, MAS adopt a decentralized architecture in which agents interact locally while achieving global coordination. This decentralized design offers notable advantages in scalability, fault tolerance, and adaptability, particularly in dynamic, partially observable, or non-stationary environments. Core attributes of MAS include autonomy, decentralization, inter-agent communication, local information access, and dynamic adaptability. These features collectively empower MAS to tackle a broad spectrum of complex, high-dimensional tasks that demand parallel sensing, distributed planning, and real-time coordination, with prominent applications in domains such as robotics, autonomous driving, and smart infrastructure.
@@ 2.3 Optimal Control
Formally, a standard OC problem can be described as follows: Given a dynamic system governed by $\dot{x}(t)=f(x(t),u(t),t)$ and an initial state $x(t_{0})=x_{0}$, the goal is to find a control input $u(t)$ that minimizes the performance index:
@@ 2.4 Reinforcement Learning
Reinforcement Learning (RL) is a fundamental branch of machine learning that focuses on enabling agents to make sequential decisions by interacting with dynamic environments [171]. Unlike supervised learning, which requires labeled data, or unsupervised learning, which uncovers hidden patterns in unlabeled data, RL is grounded in a trial-and-error process. Through continuous interaction, agents learn to select actions that maximize long-term cumulative rewards, using feedback from the environment to improve their behavior over time. Owing to its capacity for learning from interaction, RL has achieved impressive results across a broad range of domains. These include strategic game playing [93, 164], robotic control [160], industrial process optimization [130], personalized healthcare [222], autonomous driving [88], and even alignment and instruction-following in LLMs [137, 66].
@@
where the expectation is taken over trajectories generated by the policy $\pi$ and the transition dynamics $P$. While single-agent RL provides a solid foundation for autonomous decision-making, many real-world scenarios involve multiple agents acting and learning simultaneously. This motivates the extension of RL into the MARL setting, where coordination, competition, or cooperation among agents becomes essential. Specifically, in fully cooperative tasks, MARL problems are often modeled using the decentralized partially observable Markov decision process (Dec-POMDP) framework, as shown in Figure 5(b). A Dec-POMDP extends the MDP by introducing multiple agents and partial observability. Formally, it is defined as a tuple $\mathcal{M}=\langle I,\mathcal{S},{\mathcal{A}_{i}},P,R,{\Omega_{i}},O,n,\gamma\rangle$, where $I$ is the set of $n$ agents. Each agent $i$ selects actions from its own action space $\mathcal{A}_{i}$, based solely on local observations from $\Omega_{i}$. The global state transitions according to the joint action $\bm{a}$ via $P(s^{\prime}|s,\bm{a})$, and all agents receive a shared reward $R(s,\bm{a})$. Observations are generated through the function $O(s^{\prime},\bm{a},\bm{o})$, providing each agent with private, potentially incomplete information. The central challenge lies in learning a set of decentralized policies $[\pi_{1},\ldots,\pi_{n}]$, where each agent must act optimally using only local information, yet their joint behavior maximizes the team’s cumulative reward.
A wide range of algorithmic strategies has been developed to address both single-agent and multi-agent RL problems. Among these, two principal categories stand out: value-based methods and policy-based methods. Value-based methods, such as the deep Q-network (DQN) [125], focus on estimating action-value functions, enabling agents to act greedily with respect to predicted future rewards. In contrast, policy-based methods, such as DDPG [102] and PPO [153], directly optimize parameterized policies. These methods often incorporate value functions to guide learning and are especially suited to environments with high-dimensional or continuous action spaces. Both paradigms have been extended to the multi-agent setting through distinct architectural innovations. Value decomposition methods, such as QMIX [145], learn a global value function that can be decomposed into individual agent utilities. This structure supports decentralized execution while enabling centralized training. On the other hand, policy-based MARL approaches, including MADDPG [112] and MAPPO [223], adopt the CTDE framework. These methods employ a shared global critic during training to provide stabilized learning signals and support effective credit assignment among agents, thereby enhancing coordination in cooperative tasks.
@@ 2.5 Hierarchical Learning
The core process of hierarchical learning typically involves two stages: low-level policy learning and high-level policy learning. Low-level policy learning is aimed at mastering basic subtasks and is often implemented using traditional control methods such as MPC [163, 47, 42], or through end-to-end RL [218, 97]. In contrast, high-level policy learning is responsible for orchestrating the learned low-level skills to accomplish more complex goals. This is commonly achieved through RL or via planning with LLMs. In RL–based high-level policy learning, the action space is defined as the set of learned low-level policies. A parameterized high-level policy is then trained using reward signals from the environment to select and sequence these low-level skills effectively [218, 97]. In contrast, LLM-based high-level policy learning typically involves providing the task objective and the available set of low-level policies as input to the LLM, which then directly generates a structured plan by composing and invoking the appropriate low-level skills to accomplish the task [78].
@@ 2.7 Generative Models
Generative models constitute a foundational paradigm in machine learning, with the central objective of capturing the underlying distribution of training data to generate novel samples that exhibit similar characteristics. These models have found widespread applications across domains such as vision, language, and multi-modal learning. In recent years, the emergence of large-scale generative models, such as LLMs and VLMs, has significantly advanced the field. Their success is largely attributed to strong generalization capabilities, the availability of massive datasets, and scalable architectures. At the core of these models lie several key architectural frameworks, including Transformers [181], diffusion models [70], and more recently, state space models (SSMs) such as Mamba [60].
Among these, the Transformer architecture has played a pivotal role in revolutionizing sequence modeling. Originally proposed for machine translation, Transformers eliminate the need for recurrence or convolution by introducing an attention-based mechanism that enables each element in a sequence to directly attend to every other element (cf. Figure 8(a)). This design facilitates efficient parallel computation and allows the model to capture global contextual dependencies. The core attention mechanism is mathematically defined as:
@@
While Transformers and diffusion models have achieved remarkable success, both architectures encounter limitations when scaling to long sequences or complex generative processes. To address these challenges, the Mamba architecture introduces a novel approach based on continuous-time SSMs, offering linear-time complexity and improved efficiency in long-sequence processing (cf. Figure 8(c)). Mamba represents sequential data using the following differential equations:
@@ 3.1 Classic Control and Planning
To accomplish long-horizon tasks in the physical world, embodied agents must effectively plan and control their motions, generating trajectories grounded in both perceptual and task-specific inputs. Classic planning methods can be broadly categorized into constraint-based, sampling-based and optimization-based approaches [252]. Constraint-based methods encode task objectives and environmental conditions as logical constraints, translating the planning domain into symbolic representations, and employ constraint-solving techniques, such as symbolic search, to identify feasible solutions [113, 34]. However, these methods typically focus on feasibility rather than solution quality, often neglecting optimality. Additionally, when perceptual inputs are high-dimensional or structurally complex, the complexity of constraint solving increases substantially. Sampling-based methods address these limitations by incrementally exploring the feasible solution space through random sampling techniques, such as rapidly-exploring random trees (RRT) and its variants, progressively building a tree or graph structure to discover viable motion trajectories [94, 225, 136].
@@
To further optimize motion planning within the feasible domain, optimization-based methods explicitly model task objectives and performance metrics as objective functions, while representing feasibility conditions as hard constraints. Advanced optimization techniques are then utilized to search for optimal solutions within the constrained solution space. Representative optimization-based methods include polynomial trajectory planning [124], MPC [91], and OC [48]. Among these, polynomial trajectory planning generally exhibits relatively low computational cost, ensuring real-time responsiveness and smooth trajectory generation, making it particularly suitable for autonomous navigation tasks in complex scenarios [220, 143, 174, 192, 50]. In contrast, MPC and OC approaches are better suited for highly constrained scenarios requiring time-optimal maneuvers, such as collision avoidance during UAV operations [48, 57, 257, 226]. For example, illustrated in Figure 9, the work [226] introduces an aggressive collision-inclusive motion planning framework that strategically utilizes collisions to improve navigation efficiency in constrained environments. This method integrates an enhanced collision-inclusive rapidly-exploring random tree (ICRRT) for generating deformable motion primitives, alongside a trajectory optimization approach guided by a receding horizon decision (RHD) strategy for autonomous collision handling. Experimental results demonstrate the effectiveness of this approach, showing a substantial reduction (approximately 58.2%) in task completion time compared to conventional collision-free methods, underscoring its considerable potential for practical applications.
@@ 3.2 Learning based Methods
Traditional control and planning methods have long been employed for constrained real-time decision-making due to their ability to produce high-precision solutions. However, these approaches are often computationally intensive, limiting scalability and responsiveness, especially in high-dimensional, nonlinear, or non-stationary systems. Moreover, they typically exhibit poor generalization to unseen scenarios or dynamically changing environments. To overcome these limitations, learning-based decision-making paradigms have gained increasing attention, aiming to deliver real-time performance with improved robustness and generalization by learning directly from interactive data. Among these, end-to-end RL has emerged as a prominent approach, as it enables direct policy optimization through interaction with the environment. Nonetheless, in tasks characterized by complex objectives or large state-action spaces, end-to-end RL suffers from low sample efficiency and prolonged training times due to the expansive exploration space. To alleviate this issue, hierarchical RL has been introduced, which decomposes complex tasks into a sequence of simpler subtasks, thereby improving learning efficiency and scalability. Despite these advances, both end-to-end and hierarchical RL approaches fundamentally depend on the availability of well-designed reward functions, which are often difficult or impractical to define in real-world settings. In such scenarios, IL provides a compelling alternative by enabling agents to acquire effective policies directly from expert demonstrations, circumventing the need for explicit reward engineering and facilitating faster deployment in practice.
@@
In sophisticated environments, OC problems often exhibit non-convex characteristics, making it challenging for classic planning and control methods to identify feasible solutions within a reasonable computational time. A natural approach to addressing this issue is to optimize the entire system in an end-to-end manner, employing RL techniques to directly map perceptual information to action decisions through neural networks. These methods optimize the policy by maximizing the expected cumulative rewards, as defined in (3), ensuring stable and high-frequency control outputs (cf. Figure 10). Consequently, end-to-end learning is widely recognized as an ultimate pathway toward embodied AI, as it removes the necessity for manual decomposition, alignment, and tuning of individual modules, enabling agents to directly translate raw sensory inputs into effective actions. Indeed, end-to-end RL has achieved impressive performance on diverse robotic platforms equipped with multi-modal inputs, including UAVs [209, 52], autonomous vehicles [53], robotic manipulators [56], and legged robots [15, 67].
@@
Unlike OC methods, learning-based approaches such as RL require extensive data collection through interaction with the environment in a trial-and-error manner, incurring high costs in real-world embodied settings. Furthermore, RL often relies heavily on meticulously designed objective functions or reward signals, which become challenging to specify accurately in complex and unstructured environments due to the necessity for flexibility and adaptability. In these scenarios, learning from expert demonstrations, commonly referred to as IL, presents a more practical alternative. For example, [81] successfully transferred a painting policy from a simulator to a real-world robotic arm using the classic IL method BC, demonstrating high-level painting and calligraphy skills, as shown in Figure 12. Beyond this, a range of more advanced IL techniques have been introduced to further enhance embodied intelligence with improved generalization and robustness, including interactive IL [133], constrained IL [156], and IRL [235].

### v2

@@ 2 Preliminaries
In this section, we present the core techniques that underpin Embodied AI, starting with a formal definition of Embodied AI itself. We also define the concept of MAS and explore a range of essential methods, including optimal control (OC), RL, MARL, hierarchical learning, IL, and generative models.
@@ 2.1 Embodied AI
Embodied AI (Figure 2) refers to a class of intelligent agents equipped with physical bodies that enable them to perceive, act upon, and adapt to the environments through continuous interaction [112]. The concept of embodied AI trace back to Alan Turing’s early proposition in the 1950s, suggesting that genuine intelligence must arise from sensory and motor experiences rather than purely symbolic computation [185]. This notion was further formalized in the 1980s through embodied cognition theory, which posits that cognition is inherently shaped by the agent’s physical form and interaction with the world [19]. In contrast to traditional AI paradigms that rely on abstract reasoning or passive learning from static datasets, embodied AI emphasizes real-world interaction as the foundation for learning and decision-making.
At a system level, embodied AI architectures are typically composed of three tightly integrated components: perception, cognition, and action. Firstly, agents use sensors to acquire real-time, context-dependent information from their surroundings. Then the sensory data is processed by cognitive modules that support reasoning, interpretation and planning. Finally, resulting decisions are translated into actions via actuators, which modify the environment and initiate new perceptual inputs. These processes form a continuous feedback loop known as the perception–cognition–action cycle [16], which enables embodied agents to adapt their behavior dynamically based on feedback. The paradigm of embodied AI mainly include three foundational attributes that govern how intelligence emerges and evolves:
  • Embodiment Embodied AI is rooted in physical agents equipped with the capabilities to perceive, move, and act in the real world. These agents come in diverse forms, including humanoid robots, quadrupeds, unmanned vehicles and drones. The physical body serves not only as the medium through which the agent interacts with its environment, but also as the structural foundation that constrains and enables its behavioral repertoire. The morphology, sensorimotor fidelity and actuation of the body jointly define the range and granularity of the agent’s interactions, shaping the scope of its situated intelligence.
@@
Recently, the advancements in generative models such as LLMs [259] have further expanded the cognitive capabilities of embodied agents. Through leveraging their powerful reasoning and generalization abilities, LLMs enable embodied agents to understand language instructions, ground semantic knowledge in physical experiences, and perform zero-shot or few-shot adaptations. These developments have accelerated the deployment of embodied AI in various real-world domains such as robotics, autonomous driving, intelligent manufacturing and healthcare [118]. Importantly, embodied AI is not merely the integration of powerful AI models with robotic platforms. It rather represents a co-evolutionary paradigm where intelligent algorithms which serve as brains, physical structures which serve as bodies and dynamic environments evolve jointly to support adaptive embodied intelligence.
@@ 2.2 Multi-Agent Systems
MAS are composed of multiple autonomous agents, each capable of perceiving its environment, making independent decisions and executing actions accordingly [38]. In contrast to centralized control paradigms, MAS adopt a decentralized architecture in which agents interact locally while achieving global coordination. This decentralized design offers notable advantages in scalability, fault tolerance, and adaptability, particularly in dynamic, partially observable, or non-stationary environments. Core attributes of MAS include autonomy, decentralization, inter-agent communication, local information access, and dynamic adaptability. These features collectively empower MAS to tackle a broad spectrum of complex, high-dimensional tasks that demand parallel sensing, distributed planning, and real-time coordination, with prominent applications in domains such as robotics, autonomous driving, and smart infrastructure.
@@ 2.3 Optimal Control
Formally, a standard OC problem can be described as follows: Given a dynamic system governed by $\dot{x}(t)=f(x(t),u(t),t)$ and an initial state $x(t_{0})=x_{0}$, the goal is to find the optimal control input $u(t)$ that minimizes the performance index:
@@ 2.4 Reinforcement Learning
Reinforcement Learning (RL) is a fundamental branch of machine learning that focuses on enabling agents to make sequential decisions by interacting with dynamic environments [177]. Unlike supervised learning, which requires labeled data, or unsupervised learning, which uncovers hidden patterns in unlabeled data, RL is grounded in a trial-and-error process. Through continuous interaction, agents learn to select actions that maximize long-term cumulative rewards, using feedback from the environment to improve their behavior over time. Owing to its capacity for learning from interaction, RL has achieved impressive results across a broad range of domains. These include strategic game playing [95, 170], robotic control [166], industrial process optimization [134], personalized healthcare [229], autonomous driving [90, 104], and even alignment and instruction-following in LLMs [141, 67].
@@
where the expectation $\mathbb{E}_{\pi}[\cdot]$ is taken over trajectories generated by the policy $\pi$ and the transition dynamics $P$. While single-agent RL provides a solid foundation for autonomous decision-making, many real-world scenarios involve multiple agents acting and learning simultaneously. This motivates the extension of RL into MARL, where coordination or competition among agents becomes essential. Specifically, in fully cooperative tasks, MARL problems are often modeled as the decentralized partially observable Markov decision process (Dec-POMDP) framework as shown in Figure 5(b). A Dec-POMDP extends the MDP by introducing multiple agents and partial observability. Formally, it is defined as a tuple $\mathcal{M}=\langle I,\mathcal{S},{\mathcal{A}_{i}},P,R,{\Omega_{i}},O,n,\gamma\rangle$, where $I$ is the set of $n$ agents. Each agent $i$ selects actions from its own action space $\mathcal{A}_{i}$, based solely on local observations from $\Omega_{i}$. The global state transitions according to the joint action $\bm{a}$ via $P(s^{\prime}|s,\bm{a})$, and all agents receive a shared reward $R(s,\bm{a})$. Observations are generated through the function $O(s^{\prime},\bm{a},\bm{o})$, providing each agent with private, potentially incomplete information. The central challenge lies in learning a set of decentralized policies $[\pi_{1},\ldots,\pi_{n}]$, where each agent must act appropriately using local information, yet their joint behavior maximizes the team’s cumulative reward.
A wide range of algorithmic strategies has been developed to address both single-agent and multi-agent RL problems [22, 265, 86]. Among these, two principal categories stand out: value-based methods and policy-based methods. Value-based methods, such as the deep Q-network (DQN) [129], focus on estimating action-value functions, enabling agents to act greedily with respect to predicted future rewards. In contrast, policy-based methods, such as DDPG [105] and PPO [159], directly optimize parameterized policies. These methods often incorporate value functions to guide learning and are especially suited to environments with high-dimensional or continuous action spaces. Both paradigms have been extended to the multi-agent settings through distinct architectural innovations. Value decomposition methods, such as QMIX [150], learn a global value function that can be decomposed into individual agent utilities. This structure supports decentralized execution while enabling centralized training. On the other hand, policy-based MARL approaches, including MADDPG [115] and MAPPO [230], adopt the CTDE framework. These methods employ a shared global critic during training to provide stabilized learning signals and support effective credit assignment among agents, thereby enhancing coordination in cooperative tasks.
@@ 2.5 Hierarchical Learning
The core process of hierarchical learning typically involves two stages: low-level skill learning and high-level policy learning. The low-level skill learning is aimed at mastering basic sub-tasks and is often implemented using classic controllers such as MPC [169, 48, 43], or through end-to-end RL [225, 99]. In contrast, the high-level policy learning is responsible for orchestrating the learned low-level skills to accomplish more complex goals. This is commonly achieved through RL or via planning with LLMs. In RL–based high-level policy learning, the action space is defined as the set of learned low-level policies. A parameterized high-level policy is then trained using reward signals from the environment to select and sequence these low-level skills effectively [225, 99]. In contrast, LLM-based high-level policy learning typically involves providing the task objective and the available set of low-level skills as input to the LLM, which then directly generates a structured plan by composing and invoking the appropriate low-level skills to accomplish the task [79].
@@ 2.7 Generative Models
Generative models constitute a foundational paradigm in machine learning, with the central objective of capturing the underlying distribution of training data to generate novel samples that exhibit similar characteristics. These models have found widespread applications across domains such as vision, language and multi-modal learning. In recent years, the emergence of large-scale generative models, such as LLMs and VLMs, has significantly advanced the field. Their success is largely attributed to strong generalization capabilities, the availability of massive datasets and scalable architectures. At the core of these models lie several key architectural frameworks, including Transformers [187], diffusion models [71] and more recently, state space models (SSMs) such as Mamba [61]. Among these, the Transformer architecture has played a pivotal role in revolutionizing sequence modeling. Originally proposed for machine translation, Transformers eliminate the need for recurrence or convolution by introducing an attention-based mechanism that enables each element in a sequence to directly attend to every other element (cf. Figure 8(a)). This design facilitates efficient parallel computation and allows the model to capture global contextual dependencies. The core attention mechanism is mathematically defined as:
@@
While Transformers and diffusion models have achieved remarkable success, they both encounter limitations when scaling to long sequences or complex generative processes. To address these challenges, the Mamba architecture introduces a novel approach based on continuous-time SSMs, offering linear-time complexity and improved efficiency in long-sequence processing (cf. Figure 8(c)). Mamba represents sequential data using the following differential equations:
@@ 3.1 Classic Control and Planning
To accomplish long-horizon tasks in the physical world, embodied agents must effectively plan and control their motions, generating trajectories grounded in both perceptual and task-specific inputs. To this end, a rich body of classic planning methods has been developed, which can be broadly grouped into constraint-based, sampling-based, and optimization-based approaches [260]. Constraint-based methods encode task objectives and environmental conditions as logical constraints, translating the planning domain into symbolic representations, and employ constraint-solving techniques, such as symbolic search, to identify feasible solutions [116, 35]. However, these methods typically focus on feasibility rather than solution quality, often neglecting optimality. Additionally, when perceptual inputs are high-dimensional or structurally complex, the difficulty of constraint solving increases substantially. Sampling-based methods address these limitations by incrementally exploring the feasible solution space through random sampling techniques, such as rapidly-exploring random trees (RRT) and its variants, progressively building a tree or graph structure to discover viable motion trajectories [96, 232, 140].
@@
To further optimize motion planning within the feasible domain, optimization-based methods explicitly model task objectives and performance metrics as objective functions, while representing feasibility conditions as hard constraints. Advanced optimization techniques are then utilized to search for optimal solutions within the constrained solution space. Representative optimization-based methods include polynomial trajectory planning [127], MPC [93] and OC [49]. Among these, polynomial trajectory planning generally exhibits relatively low computational cost, ensuring real-time responsiveness and smooth trajectory generation, making it particularly suitable for autonomous navigation tasks in complex scenarios [227, 147, 180, 199, 51]. In contrast, MPC and OC approaches are better suited for highly constrained scenarios requiring time-optimal maneuvers, such as collision avoidance during UAV operations [49, 58, 266, 233]. For example, as illustrated in Figure 9, the work [233] introduces an aggressive collision-inclusive motion planning framework that strategically utilizes collisions to improve navigation efficiency in constrained environments. This method integrates an enhanced collision-inclusive rapidly-exploring random tree (ICRRT) for generating deformable motion primitives, alongside a trajectory optimization approach guided by a receding horizon decision strategy for autonomous collision avoidance. Experimental results demonstrate a substantial reduction in task completion time compared to conventional collision-free methods, showing the effectiveness of this approach and underscoring its considerable potential for practical applications.
@@ 3.2 Learning based Methods
Classic control and planning methods have long been employed for constrained real-time decision-making due to their ability to produce high-precision solutions. However, these approaches are often computationally intensive, limiting scalability and responsiveness, especially in high-dimensional, nonlinear or non-stationary systems. Moreover, they typically exhibit poor generalization to unseen scenarios or dynamic environments. To overcome these limitations, learning-based decision-making paradigms have gained increasing attention, aiming to deliver real-time performance with improved robustness and generalization by learning directly from interactive data [195]. Among these, end-to-end RL has emerged as a prominent approach as it enables direct policy optimization through interaction with the environment. Nonetheless, in tasks with complex objectives or large state-action spaces, end-to-end RL suffers from low sample efficiency and prolonged training times due to the expansive exploration space. To alleviate this issue, hierarchical RL has been introduced, which decomposes overall tasks into a sequence of sub-tasks, thereby improving learning efficiency and scalability. Despite these advances, both end-to-end and hierarchical RL approaches fundamentally depend on the availability of well-designed reward functions, which are often difficult or impractical to define in real-world settings. In such scenarios, IL provides a compelling alternative by enabling agents to acquire effective policies directly from expert demonstrations, circumventing the need for explicit reward engineering and facilitating faster deployment in practice.
@@
In sophisticated environments, OC problems often exhibit non-convex characteristics, making it challenging for classic planning and control methods to identify feasible solutions within a reasonable computational time. A natural approach to addressing this issue is to optimize the entire system in an end-to-end manner, employing RL techniques to directly map perceptual information to action decisions through neural networks. These methods optimize the policy by maximizing the expected cumulative rewards as defined in (3), ensuring stable and high-frequency control outputs (cf. Figure 10). Consequently, end-to-end learning is widely recognized as an ultimate pathway toward embodied AI, as it removes the necessity for manual decomposition, alignment, and tuning of individual modules, enabling agents to directly translate raw sensory inputs into effective actions. Indeed, end-to-end RL has achieved impressive performance on diverse robotic platforms equipped with multi-modal inputs, including UAVs [216, 53], autonomous vehicles [54], manipulators [57] and legged robots [15, 68].
@@
Unlike OC methods, learning-based approaches such as RL require extensive data collection through interaction with the environment in a trial-and-error manner, incurring high costs in real-world embodied settings. Furthermore, RL often relies heavily on meticulously designed objective functions or reward signals, which become challenging to specify accurately in complex and unstructured environments due to the necessity for flexibility and adaptability. In these scenarios, learning from expert demonstrations, commonly referred to as IL, presents a more practical alternative. For example, [82] successfully transfers a painting policy from a simulator to a real-world robotic arm using BC, demonstrating high-level painting and calligraphy skills as shown in Figure 12. Beyond this, a range of more advanced IL techniques have been introduced to further enhance embodied intelligence with improved generalization and robustness, including interactive IL [137], constrained IL [162], and IRL [242].